OpenStudy (anonymous):

HELP! I'll give medal if you help me. Jane walked 4/5 mile on Monday. On Tuesday, she walked 1/2 of the distance she walked on Monday. How many miles did Jane walk on Tuesday?

OpenStudy (anonymous):

So essentially, we want to know what 1/2 of the distance of 4/5 of a mile is. This calls for a multiplication problem! \[\frac{ 1 }{ 2 } \times \frac{ 4 }{ 5 }\] Which is \[\frac{ 4 }{ 10 }\] Simplified, it is \[\frac{ 2 }{ 5 }\] of a mile. Hope this helps.

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):